What would you expect the nominal rate of interest to be if the real rate is 5.0 percent and the expected inflation rate is 3.0 ​percent? The nominal rate of interest would be nothing​%. ​(Round to two decimal​ places.)
Answer #1

HI,

Here real rate r = 5%

inflation i = 3%

nominal rate R=?

as per fisher equation

(1+R) = (1+i)(1+r)

1+R = (1+3%)(1+5%)

1+R = 1.03*1.05

R = 8.15%

Hence nominal rate is 8.15%

Thanks